A bag contains 2 orange marbles and 3 green marbles. A marble is picked at random from the bag then replaced. The bag then is well shaken and a second marble is taken from the same bag.
a. What is the probability that both marbles taken were orange?
b. If the first marble was not replaced,then what would be the probability that both marbles chosen were orange?

Answer by Shai(25)

given :
A bag contains two orange marbles and 3 green marbles.
A marble is randomly picked and replaced now to answer the first part
1) Probability that both were orange, since the ball picked is then replaced, you
need to be aware that the same ball can be use for calculation.
thus implying no. of orange balls that can be picked up = (2*2) and thus number
of green balls 3C0,the total possible outcomes here can be given by 5C2.
Thus Probability that both the marbles were orange(after replacement)=(2*2*3C0)/5C2= 0.4(2/5)
2)if the first marble was not replaced then we have : (2C2 * 3C0)/5C2= 0.1(1/10)
since,here once we use an orange marble, we cannot use it back,as there is no replacement.